ball of mass 0.3 kg moving with a speed of 6.0 ms-1 strikes a wall at an angle of 60° to the wall. It then rebounds at the same speed and the same angle. It is in contact with the wall for 10 ms. Calculate the impulse and the average force.

Grade:12th pass

1 Answers

Arun
6 years ago
 
F (average) = m (vf - vi)/t
V=0 m/s
V=6 m/s
Therefore by the formula,
Favg. = m (vf - vi)/t
         = 0.3 (0-6)/0.010
         = - 1.8/0.010
= - 180 N
 
Regards
Arun (askIITians forum expert)